The activation energy of a certain uncatalysed reaction at 300K is 76 kJ/mol.The activation energy is lowered by 19 kJ/mol by the use of catalyst. By what factor, the rate of catalysed reaction is lowered?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

K = Ae-E/RT                                      (E is activation energy)

rate(R)= K(a)n                     (k=rate constant & a is concentration  , n is order of reaction)                  

rate directily proportional to rate constant so

 

R1/R2 = K1/K2 = e[-E1+E2]/RT           

            =e[-76+19]/RT 

R1/R2 = e-2.28                                       (e = 2.8 approx)

R2 = 8R1